A 62.8 kg woman stands on a horizontal surface.

A) What is the volume of the woman's body (in m3) if her average density is 983 kg/m3?

I know the answer for this is 0.064 m^3.

B) What average pressure (in Pa) from her weight is exerted on the horizontal surface if her two feet have a combined area of 3.64 ✕ 10−2 m2?

______pa?

I cant figure out part b someone please help.
Answer #1

Part A. (Already Solved)

Volume = mass/density

Volume = 62.8/983 = 0.064 m^3

Part B.

We know that

Pressure = Force exerted/Area

Force exerted on ground = Weight of woman = m*g

Area = Combined area of feet = 3.64*10^-2 m^2

So,

Pressure = m*g/A

P = 62.8*9.81/(3.64*10^-2)

P = 16924.95 Pa

P = 1.69*10^4 Pa = Average Pressure exerted on ground

Please Upvote.